# Physics-informed renormalisation group flows

### v2

@@ II.1 Generating functionals with composite fields
where the superscript $\hat{\ }$ indicates the microscopic fields or operators. With 3, the existence of the generating functionals simply entails the convergence of the respective Taylor expansion at least as an asymptotic series. Note, that such a setup does not explicitly rely on a path integral representation with the measure 3. While this is a mere tautology as proving the existence of a generating functional is tantamount to proving the existence of a path integral representation, it highlights the fact that solving the flow equation is equivalent to constructing the path integral.
@@
$\Gamma_{\phi}[\phi]=\sup_{J_{\phi}}\left(\int_{\boldsymbol{x}}J^{a}_{\phi}\,\phi^{a}-\ln Z[J_{\phi}]\right)\,,\quad J_{\phi}^{a}=\frac{\delta\Gamma_{\phi}}{\delta\phi_{a}}\,.$ (9)
This leaves us with an effective action that depends on a collection of composite (mean) fields $\phi=\langle\hat{\phi}\rangle$,
@@ II.2 Generalised functional flows
where $t$ is the (negative) RG-time and $k_{\textrm{ref}}$ is some reference scale, commonly the initial scale $k_{\textrm{ref}}=\Lambda$ is chosen. We have also introduced the regulator function $R_{k}$ that encodes the details of the cutoff procedure. In the present work we concentrate on infrared momentum cutoffs, for more details see Appendix D. However, the approach is far more general as we may substitute the momentum scale $k$ by a general flow parameter $s$. This flow parameter may be a length scale or a time, or simply comprise a general reparametrisation. In this more general case, $R_{s}$ implements the respective change. The general flow uses the propagator of the composite fields,
$G_{\phi_{i}\phi_{j}}[\phi](\boldsymbol{p},\boldsymbol{q})=\left[\frac{1}{\Gamma_{\phi}^{(2)}+R_{k}}\right]_{\phi_{i}\phi_{j}}(\boldsymbol{p},\boldsymbol{q})\,,$ (13)
@@
The constant flow $\partial_{t}\ln{\cal N}_{\phi}$ on the right-hand side of 11 originates in a $k$-dependent normalisation of the path integral. In standard Wetterich flows, this constant is typically either dropped or chosen such that the value of the effective action or the effective potential at the minimum is vanishing. However, in 11 it is not a mere constant, instead it also induces a field-dependent change of $\dot{\phi}$, and hence influences the field dependence of the effective action. This is discussed later in an explicit $\phi^{4}$-example in Appendix A.
Equation 11 also contains a shift $c_{\phi}\int\phi$ in the effective action, which takes into account potential explicit symmetry breaking terms and only contains the zero mode or collective excitation of the field $\phi$. The occurrence of such a term has been observed first and discussed in detail in [15] for explicit chiral symmetry breaking in QCD.
Finally, if the composite field equals the fundamental field, $\phi=\varphi$, the terms proportional to $\dot{\phi}$ in 11 vanish and the flow equation reduces to the standard Wetterich flow [28], see also [29, 30]. In comparison to the Wetterich equation, the addition on the left-hand side of the equation considers the immediate effect of the change of field-basis on the effective action. It only has an implicit regulator dependence and could we may also used it to change the field-basis without lowering the cutoff scale $k$, hence inducing a regulator-independent flow of the field. Finally, on the right hand side of the equation we obtain a correction proportional to the regulator $R_{k}$, which corresponds to a correction of the RG flow.
@@ II.2.1 Key properties and examples
A central property of the generalised flow 11 is the regularisation of the propagator of the composite field. In the path integral representation this regularisation is introduced with an additional quadratic term of the composite field $\phi$
$\Delta S_{k}[\hat{\phi}_{k}]=\frac{1}{2}\int_{p}\hat{\phi}_{k}R_{\phi}\hat{\phi}_{k}\,,$ (15)
see [6]. This is reflected by the occurrence of the $\phi$-propagator in 11 without any transformation matrices. Importantly this direct regularisation of the composite field suppresses potential soft or even massless composite modes. If such composite soft modes are not regularised, the flow is not momentum local, for a comprehensive discussion see [4]. This leads to a worse convergence of any expansion scheme and such a setup even fails in low orders of the given expansion scheme, for a respective analysis see e.g. [31, 32].
The generalised flow equation 11 allows us to enforce field transformations. The best studied case in the literature is the introduction of dynamical composite degrees of freedom. This is well-tested and quantitatively successful in theories with bilinear fermionic composites such as QCD with the composite pions and $\sigma$-mode, as well as in ultracold atoms with composite atomic dimers and trimer fields. Then, the physics of the respective four-Fermi scattering vertex $\lambda_{\psi}$ is stored in the dynamics of the emergent composite. Practically this is implemented with the constraint,
$\dot{\phi}_{k}[\phi]:\quad\partial_{t}\lambda_{\psi}\stackrel{{\scriptstyle!}}{{=}}0\,,$ (16)
which fixes the flow of the transformation $\dot{\phi}_{k}[\phi]$, see [7, 9, 10, 11, 12, 13, 14, 15, 16]. A first application of non-linear transformations can be found in [20] within an O($N$)-theory. There, the composite field $\phi_{k}$ interpolates between the Cartesian field basis $(\varphi_{1},...,\varphi_{n})$ in the symmetric phase and the polar basis in the broken phase (Goldstonisation). The specific approach and results from [20] have also been taken over in [21]. More recently, the generalised flow has been used to eliminate the wave function $Z_{\phi}[\phi](p)$, enforcing $Z_{\phi}\equiv 1$. In terms of the flow this entails
$\partial_{t}Z_{\phi}[\phi](p)\stackrel{{\scriptstyle!}}{{=}}0\,\quad\longrightarrow\quad\dot{\phi}_{k}[\phi]\,,$ (17)
which has been implemented within the first order of the derivative expansion with $Z_{\phi}=Z_{\phi}(\phi)$ [26, 2], for an application in gravity see [33].
@@ II.2.2 Generalised effective action
We close this discussion of the generalised effective action with a remark on microscopic and macroscopic transformations as introduced in [34]. There, macroscopic transformations are done on the level of the effective action and hence re-parametrise the effective action of the fundamental field, leading to $\Gamma_{\varphi}\left[\varphi[\phi]\right]$. In turn, microscopic transformations are done on the level of the fields $\hat{\phi}$ which are also coupled to currents. The latter source terms change the effective action, leading to $\Gamma_{\phi}[\phi]$. Only on the equations of motions the two actions are identical as the current vanish. In [34] the reconstruction problem for the effective action $\Gamma_{\varphi}[\varphi]$ as well as for observables from the effective action $\Gamma_{\phi}[\phi]$ was raised. We will discuss comprehensively the computation of observables from $\Gamma_{\phi}[\phi]$ as well as the general reconstruction in Section III.3. In our opinion, the change or rather generalisation of the effective action, that is inherent to the generalised flow equations, is not a deficiency but rather is key to its potential, which is unlocked in Section III.
@@ II.3 Existence of flowing field transformations
for the sake of simplicity. Equation 18 is approached in the presence of the regulator $R_{\phi,\Lambda}=R_{\varphi,\Lambda}$ that suppresses all quantum fluctuations, and the couplings and wave functions in the classical action possibly dependent on $\Lambda$. For example, in three dimensions the mass squared scales linearly with $\Lambda$, while the self-coupling and the wave function have a finite UV limit. Note that 18 implies that the generating functional of the theory admits the path integral representation 6. This explains the notion flat field: It describes the fact that the path integral measure is flat (translation invariant) and the suppression of all quantum fluctuations leads to the initial effective action 18. If, on top of this, the initial effective action has canonical dispersions, the quantum field $\hat{\varphi}$ obeys canonical commutation relations in the absence of ultraviolet anomalous dimensions.
@@ II.3.1 Local existence of $\dot{\phi}_{k}[\phi]$
$\Delta S_{k}[\hat{\phi}_{k}]=\frac{1}{2}\sum_{a}\int_{\boldsymbol{p},\boldsymbol{q}}\hat{\phi}_{k}^{\dagger}(\boldsymbol{q})\,R_{k}(\boldsymbol{q},\boldsymbol{p})\,\hat{\phi}_{k}(\boldsymbol{p})\,,$ (26)
where the regulator function $R_{k}(\boldsymbol{q},\boldsymbol{p})$ is typically diagonal in the momenta and suppresses the propagation of modes with momenta $p_{i}^{2}\lesssim k^{2}$. We emphasise that 26 is also stabilising expansion schemes as it guarantees that the effective action is formulated in degrees of freedom, whose flows ’live’ on the same momentum shell. Amongst other stabilising properties, this optimises momentum expansions. In turn, the restriction of the regulator terms to that of the fundamental fields leads to emergent non-localities during the flow. This is most obvious in the case where the composite fields are massless. Notably this happens if the composite fields describe dynamical Goldstone modes such as in QCD with dynamical chiral symmetry breaking and the composite pion fields.
@@ III Physics-informed RG flows
Importantly, the PIRG setup stores a significant amount of information in the definition of the field, which raises the issue of reconstructing the information from the infinitesimal transformation $\dot{\phi}$ which appears in the flow equation. In Section III.3 we specifically discuss the reconstruction of correlation functions of the fundamental field $\hat{\varphi}$.
@@ III.1 Target action flows
$\partial_{t}\Gamma[\phi]-\partial_{t}\ln{\cal N}_{\phi}\stackrel{{\scriptstyle!}}{{=}}\partial_{t}\Gamma_{T}[\phi]\,,$ (29)
@@
with the full propagator of the target action $G_{T}[\phi]$. In comparison to 11, the target action flow 30 has been rearranged such that it can be read as a differential equation for the flowing fields $\dot{\phi}[\phi]$. This facilitates the discussion of general properties and constraints of the flowing fields. Indeed, 9 and 30 already lead to a first restriction on, or rather a specification of the pair 27 of, target space actions and corresponding fields $\phi$. The latter comes with a set ${\cal S}_{\phi}$ of allowed values of $\phi$ and one should consider the pair $\{\phi,{\cal S}_{\phi}\}$: By construction as a Legendre transform, only convex functionals $\Gamma_{T}[\phi]$ are allowed for $\phi\in{\cal S}_{\phi}$. As in the Wetterich equation this convexity constraint is carried by the propagator $G[\phi]$ in 30, which developed a singularity if the convexity of the full two-point function including the regulator is lost. This can be formulated as
($(\{\phi,{\cal S}_{\phi}\},\Gamma_{T})$
$\in\Bigl{\{}(\phi,\Gamma_{\phi}[\phi])\,\Big{|}\,\Gamma^{(2)}_{\phi,k}[\phi]+R^{\ }_{\phi,k}\geq 0\quad\forall\,\phi\in{\cal S}_{\phi}\Bigr{\}}\,.$ (31)
For example, for a real scalar composite field, the set ${\cal S}_{\phi}$ is a subset of the real numbers, $\phi\in\mathbbm{R}$. Then, for target actions with a convex effective action or rather $\Gamma^{(2)}_{\phi,k}[\phi]+R^{\ }_{\phi,k}\geq 0$ for all values of the field, the set ${\cal S}_{\phi}$ is simply given by the real numbers, ${\cal S}_{\phi}=\mathbbm{R}$. In turn, for $\Gamma^{(2)}_{\phi,k}[\phi]+R^{\ }_{\phi,k}<0$ for $\phi\in{\cal D}_{\phi}$, at least this subset of the real numbers has to be excluded from ${\cal S}_{\phi}\subset\mathbbm{R}\backslash{\cal D}_{\phi}$
@@
which we discuss further in Section III.3. We emphasise that within this fully factorised form the task of solving the original flow equation, which is the task of solving a functional PDE, is reduced to simply integrating the solution of a functional ODE. Evidently, this is a qualitative reduction of the numerical complexity. We can again pick up the example of the large N limit in O(N) theories, which is solved (in LPA) with the method of characteristics: Roughly speaking, the PIRG allows us to transfer the method of characteristic to general theories, with the additional restriction 68. This translates into picking the classical action as target action, which is discussed further in Section V.1.

### v3

@@ II.1 Generating functionals with composite fields
where the hat indicates the microscopic fields or operators. With 3, the existence of the generating functionals simply entails the convergence of the respective Taylor expansion at least as an asymptotic series. Note, that such a setup does not explicitly rely on a path integral representation with the measure 3. While this is a mere tautology as proving the existence of a generating functional is tantamount to proving the existence of a path integral representation, it highlights the fact that solving the flow equation is equivalent to constructing the path integral.
@@
$\Gamma_{\phi}[\phi]=\sup_{J_{\phi}}\left(\int_{\boldsymbol{x}}J^{a}_{\phi}\,\phi^{a}-\ln Z_{\phi}[J_{\phi}]\right)\,,\quad J_{\phi}^{a}=\frac{\delta\Gamma_{\phi}}{\delta\phi_{a}}\,.$ (9)
As for the generating functional $Z_{\phi}$, the subscript ϕ used in the effective action $\Gamma_{\phi}$ indicates that the Legendre transform was taken with respect to the current $J_{\phi}$. Hence, its argument is the mean field $\phi$. This leaves us with an effective action that depends on a collection of composite (mean) fields $\phi=\langle\hat{\phi}\rangle$,
@@ II.2 Generalised functional flows
where $t$ is the (negative) RG-time and $k_{\textrm{ref}}$ is some reference scale, commonly the initial scale $k_{\textrm{ref}}=\Lambda$ is chosen. In 11 we have dropped all explicit reference to the cutoff scale $k$ for the sake of readability. We have also introduced the regulator function $R_{\phi,k}$ that encodes the details of the cutoff procedure. The subscript ϕ indicates that the regulator couples to the composite fields, as discussed later around 15. In the present work we concentrate on infrared momentum cutoffs, for more details see Appendix D. However, the approach is far more general as we may substitute the momentum scale $k$ by a general flow parameter $s$. This flow parameter may be a length scale or a time, or simply comprise a general reparametrisation. In this more general case, $R_{\phi,s}$ implements the respective change. The general flow uses the propagator of the composite fields,
$G_{\phi_{i}\phi_{j}}[\phi](\boldsymbol{p},\boldsymbol{q})=\left[\frac{1}{\Gamma_{\phi}^{(2)}+R_{\phi}}\right]_{\phi_{i}\phi_{j}}(\boldsymbol{p},\boldsymbol{q})\,$, (13)
@@
[…]
The shift $c_{\phi}\int\phi$ in 11 occurs in the presence of linear (in $\phi$) explicit symmetry breaking terms in the classical action. It only affects the zero mode or collective excitation of the field $\phi$. Such a breaking can be absorbed in a shift of the current $J_{\phi}$, and the combination $\Gamma_{\phi}-c_{\phi}\int\phi$ is symmetric: This accommodates the fact that linear breaking terms do not enter the quantum dynamics of the effective action and hence the flow equation is that of the symmetric theory. The occurrence and necessity of such a term has been first discussed in detail in [15] for explicit chiral symmetry breaking in QCD. There, the respective linear terms in the fermionic bilinear composite $\phi\propto\bar{q}q$ due to the non-vanishing current quark mass term proportional to $\int\bar{q}q$ in the fundamental QCD action. More details can be found in Section II in [15]. In the present work we incorporated this term in 11 only for the sake of completeness. All examples considered here are fully O(N)-symmetric scalar theories, i.e. $c_{\sigma}=0$, and we shall consider the general case with $c_{\sigma}\neq 0$ in future work.
Finally, if the composite field equals the fundamental field, $\phi=\varphi$, the terms proportional to $\dot{\phi}$ in 11 vanish and the flow equation reduces to the standard Wetterich flow [28], see also [29, 30]. In comparison to the Wetterich equation, the addition on the left-hand side of the equation considers the immediate effect of the change of field-basis on the effective action. It only has an implicit regulator dependence and could also be used to change the field-basis without lowering the cutoff scale $k$, hence inducing a regulator-independent flow of the field. Finally, on the right hand side of the equation we obtain a correction proportional to the regulator $R_{\phi}$, which corresponds to a correction of the RG flow.
@@ II.2.1 Key properties and examples
A central property of the generalised flow 11 is the regularisation of the propagator of the composite field. In the path integral representation this regularisation is introduced with an additional quadratic term of the composite field $\phi$,
$\Delta S_{k}[\hat{\phi}_{k}]=\frac{1}{2}\int\hat{\phi}_{k}\,R_{\phi,k}\,\hat{\phi}_{k}\,$, (15)
see [6]. In 15 we concentrate on the quadratic nature in $\hat{\phi}$, a more detailed form is provided later in 26. The $\hat{\phi}^{2}$-dependence is reflected by the occurrence of the $\phi$-propagator in 11 without any transformation matrices. Importantly this direct regularisation of the composite field suppresses potential soft or even massless composite modes. If such composite soft modes are not regularised, the flow is not momentum local, for a comprehensive discussion see [4]. This leads to a worse convergence of any expansion scheme and such a setup even fails in low orders of the given expansion scheme, for a respective analysis see e.g. [31, 32].
The generalised flow equation 11 allows us to enforce field transformations. The best-studied case in the literature is the introduction of dynamical composite degrees of freedom. This is well-tested and quantitatively successful in theories with bilinear fermionic composites such as QCD with the composite pions and $\sigma$-mode, as well as in ultracold atoms with composite atomic dimers and trimer fields. Then, the physics of the respective four-Fermi scattering vertex $\lambda_{\psi}$ is stored in the dynamics of the emergent composite. Practically this is implemented with the constraint,
$\dot{\phi}_{k}[\phi]:\quad\partial_{t}\lambda_{\psi}\stackrel{{\scriptstyle!}}{{=}}0\,.$ (16)
This fixes the flow of the transformation $\dot{\phi}_{k}[\phi]$, [7, 9, 10, 11, 12, 13, 14, 15, 16].
A first application with non-linear transformations can be found in [20] within an O(N)-theory. There, the composite field $\phi_{k}$ interpolates between the Cartesian field basis $(\varphi_{1},...,\varphi_{n})$ in the symmetric phase and the polar basis in the broken phase (Goldstonisation). The specific approach and results from [20] have also been taken over in [21]. More recently, the generalised flow has been used to eliminate the wave function $Z_{\phi}[\phi](p)$, enforcing $Z_{\phi}\equiv 1$. In terms of the flow this entails, [26],
$\partial_{t}Z_{\phi}[\phi](p)\stackrel{{\scriptstyle!}}{{=}}0\,\quad\longrightarrow\quad\dot{\phi}_{k}[\phi]\,.$ (17)
This has been implemented within the first order of the derivative expansion with $Z_{\phi}=Z_{\phi}(\phi)$ [26, 2, 33], for applications in gravity see [34, 35, 36, 37].
Both non-linear examples, the Goldstonisation in [20, 21] and the removal of $Z_{\phi}[\phi]$ in 17 can be embedded in a ground state expansion, see [2]. In such an expansion, the effective action $\Gamma_{\phi}$ is expanded about the (on-shell) ground state of the theory: this entails ’classical’ dispersions as well as a parametrisation in the dynamical degrees of freedom such as the Goldstone modes in the broken phase. We emphasise that this expansion scheme also incorporates the emergent bilinear composites discussed around 16 as dynamical infrared degrees of freedom.
@@ II.2.2 Generalised effective action
We close this discussion of the generalised effective action with a remark on microscopic and macroscopic transformations as introduced in [38]. There, macroscopic transformations are done on the level of the effective action and hence re-parametrise the effective action of the fundamental field, leading to $\Gamma_{\varphi}\left[\varphi[\phi]\right]$. In turn, microscopic transformations are done on the level of the fields $\hat{\phi}$ which are also coupled to currents. The latter source terms change the effective action, leading to $\Gamma_{\phi}[\phi]$. Only on the equations of motions the two actions are identical as the currents vanish. In [38] the reconstruction problem for the effective action $\Gamma_{\varphi}[\varphi]$ as well as for observables from the effective action $\Gamma_{\phi}[\phi]$ was raised. We will discuss comprehensively the computation of observables from $\Gamma_{\phi}[\phi]$ as well as the general reconstruction in Section III.3. In our opinion, the change or rather generalisation of the effective action, that is inherent to the generalised flow equations, is not a deficiency but rather is key to its potential, which is unlocked in Section III.
@@ II.3 Existence of flowing field transformations
for the sake of simplicity. Equation 18 is approached in the presence of the regulator $R_{\Lambda}$ that suppresses all quantum fluctuations in $\phi$ and $\varphi$, and the couplings and wave functions in the classical action possibly depend on $\Lambda$. For example, in three dimensions the RG flow of the mass squared scales linearly with $\Lambda$, while the self-coupling and the wave function have a finite UV limit. Note that 18 implies that the generating functional of the theory admits the path integral representation 6. This explains the notion flat field: It describes the fact that the path integral measure is flat (translation invariant) and the suppression of all quantum fluctuations leads to the initial effective action 18. If, on top of this, the initial effective action has canonical dispersions, the quantum field $\hat{\varphi}$ obeys canonical commutation relations in the absence of ultraviolet anomalous dimensions.
@@ II.3.1 Local existence of $\dot{\phi}_{k}[\phi]$
$\Delta S_{k}[\hat{\phi}_{k}]=\frac{1}{2}\sum_{ab}\int_{\boldsymbol{p},\boldsymbol{q}}\hat{\phi}_{a,k}^{\dagger}(\boldsymbol{q})\,R^{ab}_{\phi,k}(\boldsymbol{q},\boldsymbol{p})\,\hat{\phi}_{b,k}(\boldsymbol{p})\,.$ (26)
The indices a,b{a,b} in 26 sum over different composite field species, and the regulator function $R_{\phi}(\boldsymbol{q},\boldsymbol{p})$ is typically diagonal in the momenta and suppresses the propagation of modes with momenta $p_{i}^{2}\lesssim k^{2}$. We emphasise that 26 is also stabilising expansion schemes as it guarantees that the effective action is formulated in degrees of freedom, whose flows ’live’ on the same momentum shell. Amongst other stabilising properties, this optimises momentum expansions. In turn, the restriction of the regulator terms to that of the fundamental fields leads to emergent non-localities during the flow. This is most obvious in the case where the composite fields are massless. Notably this happens if the composite fields describe dynamical Goldstone modes such as in QCD with dynamical chiral symmetry breaking and the composite pion fields.
@@ III Physics-informed RG flows
Importantly, the PIRG setup stores a significant amount of information in the definition of the field. This raises the issue of reconstructing the information from the infinitesimal transformation $\dot{\phi}$ which appears in the flow equation. In Section III.3 we specifically discuss the reconstruction of correlation functions of the fundamental field $\hat{\varphi}$.
@@ III.1 Target action flows
[…]
@@
with the full propagator of the target action $G_{T}[\phi]$. In comparison to 11, the target action flow 30 has been rearranged such that it can be read as a differential equation for the flowing fields $\dot{\phi}[\phi]$. This facilitates the discussion of general properties and constraints of the flowing fields. Indeed, 9 and 30 already lead to a first restriction on, or rather a specification of the pair 27 of, target space actions and corresponding fields $\phi$. The latter comes with a set ${\cal S}_{\phi}$ of allowed values of $\phi$ and one should consider the pair $\{\phi,{\cal S}_{\phi}\}$: $\Gamma_{T}+\Delta S_{k}$ is a Legendre transform, and hence is convex. This entails that $\Gamma^{(2)}_{T}[\phi]+\Delta S^{(2)}_{k}[\phi]\geq 0$ for $\phi\in{\cal S}_{\phi}$. As in the Wetterich equation this convexity constraint is carried by the propagator $G[\phi]$ in 30, which developed a singularity if the convexity of the full two-point function including the regulator is lost. This can be formulated as
$(\{\phi,{\cal S}_{\phi}\},\Gamma_{T})$
$\hskip-48.36958pt\in\Bigl{\{}(\phi,\Gamma_{\phi}[\phi])\,\Big{|}\,\Gamma^{(2)}_{\phi,k}[\phi]+R^{\ }_{\phi,k}\geq 0\quad\forall\,\phi\in{\cal S}_{\phi}\Bigr{\}}\,.$ (31)
For example, for a real scalar composite field, the set ${\cal S}_{\phi}$ is a subset of the real numbers, $\phi\in\mathbbm{R}$. Then, for target actions with a convex effective action or rather $\Gamma^{(2)}_{\phi,k}[\phi]+R^{\ }_{\phi,k}\geq 0$ for all values of the field, the set ${\cal S}_{\phi}$ is simply given by the real numbers, ${\cal S}_{\phi}=\mathbbm{R}$. In turn, for $\Gamma^{(2)}_{\phi,k}[\phi]+R^{\ }_{\phi,k}<0$ for $\phi\in{\cal D}_{\phi}$, at least this subset of the real numbers has to be excluded from ${\cal S}_{\phi}\subset\mathbbm{R}\backslash{\cal D}_{\phi}$ .
@@
which we discuss further in Section III.3. We emphasise that within this fully factorised form the task of solving the original flow equation, which is the task of solving a functional PDE, is reduced to simply integrating the solution of a functional ODE. Evidently, this is a qualitative reduction of the numerical complexity. We can again pick up the example of the large N limit in O(N) theories, which is solved (in LPA) with the method of characteristics: Roughly speaking, the PIRG allows us to transfer the method of characteristics to general theories. This comes at the price of some additional subtleties that will be discussed in Section III.3. The method of characteristics translates into picking the classical action as target action, which we consider in Section V.1.